# Optimal Clustering with Dependent Costs in Bayesian Networks

### v4

@@ 2.3.3 Dynamic Programming
We use dynamic programming to find optimal cluster mappings. Generally, dynamic programming is formulated as the optimisation of an objective function (or cost function) $G$ over a series of decisions or actions $u_{k}$ for $k=0,\ldots,K-1$ where the optimal cost $G^{\ast}$ over an interval $[k,K]$ is (Bellman and Dreyfus, 1962):
@@
where $U_{(k+1,K-1)}^{\ast}$ is the optimal sequence of actions $u_{k+1},\ldots,u_{K-1}$ for the interval $[k+1,K-1]$, $\mathcal{A}_{k}$ is the set of possible actions at stage $k$, $x$ is the state at stage $K$ as a consequence of optimal choice of $u_{k}$ at state $x_{k}$, and $\inf$ is the infimum operator which returns the argument that gives the minimum expression. Intuitively, given the current state $x_{k}$ and an optimal sequence of actions at stages $k+1$ to $K-1$ leading to state $x=x_{K}$, we choose the action $u_{k}$ to take at the current stage to minimise the overall cost in going from stage $k$ to $K$. This formulation is the most flexible as the cost function can vary (i.e time-varying in the literature, but we refer to as stage varying to avoid confusion with time in a DBN) across intervals $[k,K]$, and does not have to be additive. This is useful as there can be multiple clusters in a given layer of the DAG to cluster (Section 2.4). However, computational complexity grows in NP time with graph size and with the number of proposals $|\mathcal{A}_{k}|$ at each stage (LaValle, 2006).
@@
where $f(\boldsymbol{x},\boldsymbol{u})$ is a sequence of states and actions, starting at stage $k=0$, resulting in state $x$ at stage $k$ with optimal cost $G^{\ast}_{k}$, and $c$ is the transition cost of enacting action $u$ at state $x$. An additive and time-invariant system greatly reduces computational complexity. Optimisation algorithms are built around the objective function, and Equation (7) is the basis for many heuristic search techniques such as A*. A heuristic estimate of the remaining cost over stages $[k+1,K]$ is used in A* to prioritise the search and potentially reduce computation time, such as for BN learning (Lu et al., 2021).
@@ 2.5 Objective Function and Heuristic
Substituting into the general objective function of Bellman and Dreyfus (1962) (6), the objective function over layers $[l,L]$ is:
@@
Equation (10) provides the basis for our algorithm. However, in algorithms like A* (LaValle, 2006), potential actions (i.e. cluster mappings) are evaluated prospectively using the objective function at each iteration. This rapidly becomes infeasible in our cluster mapping context due to the combinatorial explosion of mappings over multiple layers $u(\boldsymbol{X}_{l+1,\ldots,L})$. One approach to overcome this challenge is to instead apply our objective function (10) retrospectively after potentially multiple iterations of the algorithm, to prune dominated branches (Section 3.2 and Lemma 3.2).
@@ 2.6 Search Space Size
As might be expected, the search space grows exponentially with layers, and the out-degree of the DAG can also significantly grow the size of the search space (Buluç et al., 2016). In- and out-degree refer to the number of parent or child nodes, respectively (Buluç et al., 2016). For instance, the case study DBN (Wu et al., 2018), with small world connectivity, has just 25 nodes but a search space of 9.9 billion mapping combinations. This study and corresponding DCMAP solution are described in Section 4.
@@ 3.1.2 Pop Queue or Activate Branch
Potentially, $Q$ could be treated as a heap to maximise computational efficiency (LaValle, 2006). We use $Pop()$ to refer to removing the lowest cost queue entry, choosing between equicost entries arbitrarily (Component Pop Queue or Activate Branch). If there are no eligible entries on the queue, i.e. no branch has $\text{Br}_{\text{pl}}\geq l$, then a branch is chosen for activation. This makes queue entries on that branch potentially eligible. Activation randomly switches between using $Peek()$ to return, but not remove, the lowest cost entry, or $Peek^{\prime}()$ to return an entry at the lowest layer on the queue, choosing arbitrarily between entries at the same layer. This facilitates switching between best-first and breadth-first search (Theorem 3.1). After activation, DCMAP tries again to pop the queue if it is not empty; the algorithm only proceeds to the next step (Component Prune Linked Dominated Branches) if a queue entry was successfully popped.
@@ 3.2 Theoretical Analysis
This section proves analytically that DCMAP is optimal in the sense that it finds all least cost cluster mappings, and does so such that an equally informed algorithm cannot do so in fewer iterations. This is achieved through conditional optimality (derived in Section 3.2.1), the exploration of all possible contiguous cluster mappings through branching (Section 3.2.2), and combined to show the optimality of DCMAP through evaluating non-dominated branches, i.e. potential cluster mapping solutions, at the earliest opportunity (Section 3.2.3). An additional property is derived for super additivity of cost functions for a heuristic to further speed up the search (Supporting Information A).
@@ 3.2.3 Optimality
Therefore, it is not possible to find, using an equally informed algorithm, all the least cost cluster mappings in fewer iterations than DCMAP in the absence of branch activation. Examples of alternate algorithms for cluster mapping include tree search algorithms like breadth first or depth first search (LaValle, 2006). However, the number of iterations needed to find the first optimal solution is also important.
@@ 4 Case Study
We use as a case study the complex systems DBN model for seagrass resilience, described briefly in Section 1. This model, which has 25 discrete nodes per time-slice (Fig. 3) and 96 time slices (Wu et al., 2018), captures biological, ecological and environmental dynamics and their interactions to predict the impact on resilience of a series of dredging stressors and its cumulative effects over time. It demonstrates small world or power law connectivity (Table 3 and 4) with a few nodes having comparatively many parent or child nodes, and the majority having few (Watts and Strogatz, 1998); most nodes had two to four states, with a few having ten or more states. Wu et al. (2018) incorporated expert knowledge and multiple data sources (e.g. experimental and monitoring data) to inform the model; however, the computation time prohibited the use of Bayesian inference, which enables better data integration and estimation of uncertainty.
@@
Figure 3: Overall seagrass DBN model (Wu et al., 2018) (Fig. 3a). Nodes are ovals and arrows denote conditional dependence between a parent and child node in the same time-slice. Where an arrow is labelled with a $1$, the child node is in the next time-slice. Rounded rectangles denote subnetworks. Yellow nodes relate to loss and recovery in shoot density, purple nodes to recovery, green nodes to resistance, blue nodes to environmental factors and pink for all other nodes. The environment subnetwork is not shown as it is a disjoint subnetwork of the blue nodes, which are effectively inputs to the model.
@@ 4.1 Computation Cost as an Objective Function
Without loss of generality, consider a discrete BN. We formulate a backwards probability measure $P^{b}_{kl}$ for computing inference akin to the rearrangement of Equation (3), which forms the basis of inference techniques such as variable elimination (Daly et al., 2011).
@@
Define $\mathcal{F}()$ as a normalised count of the number of clock cycles associated with performing all the multiplications and additions, where each multiplication is assumed to have a normalised count of $1$ cycle, each addition $0.6$ and each division (i.e. ratio) $3$, based on Intel Skylake CPU latencies (Fog, 2022). Hence,
@@ 5 Discussion
Clustering has been applied in a variety of approaches to improve the computational feasibility of BN learning, whether based on constraints, scoring (i.e. using a cost function) or both (Huang and Zhou, 2022). In this context, the vast majority of existing literature has focused on cost functions based on inclusion or exclusion of edges. Cost functions that vary with the mapping of nodes to different clusters are currently understudied. Such flexibility enables explicit optimisation of a given BN and its cluster mapping which could potentially lead to better BN learning and inference outcomes. However, finding globally optimal (i.e. least cost) cluster mappings in a BN incurs additional complexity in the form of cost dependency, where the local cost of a cluster is dependent on the mapping of connected nodes and the global cost is typically the sum of local costs.
Previous BN literature has used decomposability of scoring functions, such as through additive costs, to allow for local search procedures and/or optimisation of the score function; some even use dynamic programming and graph layers (Friedman et al., 1999; Kojima et al., 2010; Lu et al., 2021). Our work builds upon the application of dynamic programming and layers in BNs to facilitate global optimisation with arbitrary non-negative cost functions and cost dependency. This latter property has not been explicitly studied in the existing BN literature. The proposed algorithm, DCMAP, is proven to find all of the least cost cluster mapping solutions through conditional optimality of layers and branching (Section 3.2.3). In addition, it is theoretically shown that an equally informed algorithm cannot find these solutions in fewer iterations. This is important as optimal clustering is NP-hard in general (Buluç et al., 2016) and cost dependency would exacerbate this complexity. Empirically, it is found that a breadth first search offers similar performance (number of iterations) to DCMAP for the simple illustrative BN (Fig. 1). However, exponential growth in clustering complexity with layers and nodes means that breadth first search is unable to find optimal solutions on the case study DBN, even with ten times more iterations than DCMAP needs to find the first optimal solution. This DBN is also not a particularly large network.
@@ 5.1 Case Study Discussion
We developed a case study of BN clustering where clusters were used as a means to compute (i.e. infer) the marginal probabilities of a complex systems seagrass DBN. By explicitly defining the cost function as the computational cost of inference given a cluster mapping, a speed-up of 22.8 and 426 times was achieved for a one and two time-slice version of the DBN, respectively. In itself, a substantial reduction of computational complexity through clustering is useful as BN inference plays a key role in prediction and model learning, and methods for that such as Markov Chain Monte Carlo (MCMC) (Gelman et al., 2013) can require thousands of iterations. MCMC can also help better quantify uncertainty for parameter and structure learning of BNs (Daly et al., 2011) and DBNs (Shafiee Kamalabad and Grzegorczyk, 2020) through the Bayesian framework. Explicit optimisation of BN clusters using DCMAP could be a key enabler for future research in BN learning and updating, especially for complex systems DBNs like our seagrass case study.
@@
Although our case study was for a discrete BN with between two and fourteen states per node with corresponding CPTs, the proposed DCMAP algorithm equally applies to discrete or continuous or hybrid BNs, DBNs, Object Oriented Bayesian Networks (OOBNs) (Koller and Pfeffer, 2013) and Markov models in general, all of which can be represented as a DAG; the key adaptation required is the development of an application-specific cost function. Note that dynamic and templated models alike can be unrolled (Wu et al., 2018) into a BN and cluster mapping performed on the result. In practice, for a $k^{th}$ order dynamic model, running cluster mapping on $k$ slices is sufficient to find optimal clusters that can then be replicated over all slices.
@@
Although we have theoretically shown that an equally informed algorithm cannot find all optimal solutions in fewer iterations, further study is needed to better understand the computational complexity of DCMAP and how it might vary with different BN graph structures and cost functions. Early results from the case study suggest a complexity of $O(n^{2})$ iterations to find the first least cost cluster mapping and $O(an^{2})$ iterations for all optimal solutions, where $n$ is the number of nodes and $a$ some positive integer. This aligns with previous findings of $O(n^{2})$ complexity for clustering in the absence of cost dependency (Zheng et al., 2020). Furthermore, better understanding of the convergence characteristics of clustering and early termination for scenarios where only the first optimal cluster mapping is needed are also valuable areas for future work. Additionally, even though proposed cluster mappings are always unique (Lemma 3.3), label switching is an area for future study to better understand the extent of the issue and strategies for mitigation if needed.
@@ 5.2 Future Work: Entropy
In light of the generalisability of DCMAP given its ability to optimise for arbitrary, non-negative and dependent cost functions, one key area of future work is in the development of objective functions for specific usage scenarios. One such possibility is entropy, which for BNs is defined generally as the sum of the entropy associated with each scenario over all possible scenarios, where a scenario might be defined in terms of one or more nodes (Scutari, 2024; Parhizkar et al., 2018). Consider a discrete BN. The entropy associated with evidence for a given scenario indexed by $k$ is:
@@
Consider a special case of clustering where specific evidence scenarios are to be applied at one or more nodes and at one or more points in time in a DBN, with the goal of minimising global entropy. Thus, each node can be assigned to one of two clusters: (cluster 1) observed using virtual evidence $\delta(X^{\prime})$ (Pearl, 1988) where each observation is independent, or (cluster 2) not observed. Note that virtual evidence provides greater flexibility in modelling uncertainty compared to hard evidence where a state is observed or not. Dependent cluster mapping is necessary to optimise over all possible combinations of observed nodes and their interdependencies, which arise from propagation of evidence for inferring marginal probabilities for each scenario $p_{ijk}$ (20). DCMAP can approximate this marginal probability $p_{ijk}$ for every node $X_{i}$ via backwards propagation with evidence to layer $l$ and no evidence from forward propagation:
@@
Potentially, optimisation of clusters with respect to entropy could contribute to new methods in BN learning and prediction, and is a promising area for future research that is enabled by DCMAP.
Appendix A Supporting Information: Super-additive Objective Functions

### v5

@@ 2.3.3 Dynamic Programming
We use dynamic programming to find optimal cluster mappings. Generally, dynamic programming is formulated as the optimisation of an objective function (or cost function) $G$ over a series of decisions or actions $u_{k}$ for $k=0,\ldots,K-1$ where the optimal cost $G^{\ast}$ over an interval $[k,K]$ is [3]:
@@
where $U_{(k+1,K-1)}^{\ast}$ is the optimal sequence of actions $u_{k+1},\ldots,u_{K-1}$ for the interval $[k+1,K-1]$, $\mathcal{A}_{k}$ is the set of possible actions at stage $k$, $x$ is the state at stage $K$ as a consequence of optimal choice of $u_{k}$ at state $x_{k}$, and $\inf$ is the infimum operator which returns the argument that gives the minimum expression. Intuitively, given the current state $x_{k}$ and an optimal sequence of actions at stages $k+1$ to $K-1$ leading to state $x=x_{K}$, we choose the action $u_{k}$ to take at the current stage to minimise the overall cost in going from stage $k$ to $K$. This formulation is the most flexible as the cost function can vary (i.e time-varying in the literature, but we refer to as stage varying to avoid confusion with time in a DBN) across intervals $[k,K]$, and does not have to be additive. This is useful as there can be multiple clusters in a given layer of the DAG to cluster (Section 2.4). However, computational complexity grows in NP time with graph size and with the number of proposals $|\mathcal{A}_{k}|$ at each stage [16].
@@
where $f(\boldsymbol{x},\boldsymbol{u})$ is a sequence of states and actions, starting at stage $k=0$, resulting in state $x$ at stage $k$ with optimal cost $G^{\ast}_{k}$, and $c$ is the transition cost of enacting action $u$ at state $x$. An additive and time-invariant system greatly reduces computational complexity. Optimisation algorithms are built around the objective function, and Equation (7) is the basis for many heuristic search techniques such as A*. A heuristic estimate of the remaining cost over stages $[k+1,K]$ is used in A* to prioritise the search and potentially reduce computation time, such as for BN learning [18].
@@ 2.5 Objective Function and Heuristic
Substituting into the general objective function of Bellman and Dreyfus [3] (6), the objective function over layers $[l,L]$ is:
@@
Equation (10) provides the basis for our algorithm. However, in algorithms like A* [16], potential actions (i.e. cluster mappings) are evaluated prospectively using the objective function at each iteration. This rapidly becomes infeasible in our cluster mapping context due to the combinatorial explosion of mappings over multiple layers $u(\boldsymbol{X}_{l+1,\ldots,L})$. One approach to overcome this challenge is to instead apply our objective function (10) retrospectively after potentially multiple iterations of the algorithm, to prune dominated branches (Section 3.2 and Lemma 3.2).
@@ 2.6 Search Space Size
As might be expected, the search space grows exponentially with layers, and the out-degree of the DAG can also significantly grow the size of the search space [4]. In- and out-degree refer to the number of parent or child nodes, respectively [4]. For instance, the case study DBN [26], with small world connectivity, has just 25 nodes but a search space of 9.9 billion mapping combinations. This study and corresponding DCMAP solution are described in Section 4.
@@ 3.1.2 Pop Queue or Activate Branch
Potentially, $Q$ could be treated as a heap to maximise computational efficiency [16]. We use $Pop()$ to refer to removing the lowest cost queue entry, choosing between equicost entries arbitrarily (Component Pop Queue or Activate Branch). If there are no eligible entries on the queue, i.e. no branch has $\text{Br}_{\text{pl}}\geq l$, then a branch is chosen for activation. This makes queue entries on that branch potentially eligible. Activation randomly switches between using $Peek()$ to return, but not remove, the lowest cost entry, or $Peek^{\prime}()$ to return an entry at the lowest layer on the queue, choosing arbitrarily between entries at the same layer. This facilitates switching between best-first and breadth-first search (Theorem 3.1). After activation, DCMAP tries again to pop the queue if it is not empty; the algorithm only proceeds to the next step (Component Prune Linked Dominated Branches) if a queue entry was successfully popped.
@@ 3.2 Theoretical Analysis
This section proves analytically that DCMAP is optimal in the sense that it finds all least cost cluster mappings, and does so such that an equally informed algorithm cannot do so in fewer iterations. This is achieved through conditional optimality (derived in Section 3.2.1), the exploration of all possible contiguous cluster mappings through branching (Section 3.2.2), and combined to show the optimality of DCMAP through evaluating non-dominated branches, i.e. potential cluster mapping solutions, at the earliest opportunity (Section 3.2.3). An additional property is derived for super additivity of cost functions for a heuristic to further speed up the search (Supplementary Information 6).
@@ 3.2.3 Optimality
Therefore, it is not possible to find, using an equally informed algorithm, all the least cost cluster mappings in fewer iterations than DCMAP in the absence of branch activation. Examples of alternate algorithms for cluster mapping include tree search algorithms like breadth first or depth first search [16]. However, the number of iterations needed to find the first optimal solution is also important.
@@ 4 Case Study
We use as a case study the complex systems DBN model for seagrass resilience, described briefly in Section 1. This model, which has 25 discrete nodes per time-slice (Fig. 3) and 96 time slices [26], captures biological, ecological and environmental dynamics and their interactions to predict the impact on resilience of a series of dredging stressors and its cumulative effects over time. It demonstrates small world or power law connectivity (Table 3 and 4) with a few nodes having comparatively many parent or child nodes, and the majority having few [25]; most nodes had two to four states, with a few having ten or more states. [26] incorporated expert knowledge and multiple data sources (e.g. experimental and monitoring data) to inform the model; however, the computation time prohibited the use of Bayesian inference, which enables better data integration and estimation of uncertainty.
@@
Figure 3: Overall seagrass DBN model [26] (Fig. 3a). Nodes are ovals and arrows denote conditional dependence between a parent and child node in the same time-slice. Where an arrow is labelled with a $1$, the child node is in the next time-slice. Rounded rectangles denote subnetworks. Yellow nodes relate to loss and recovery in shoot density, purple nodes to recovery, green nodes to resistance, blue nodes to environmental factors and pink for all other nodes. The environment subnetwork is not shown as it is a disjoint subnetwork of the blue nodes, which are effectively inputs to the model.
@@ 4.1 Computation Cost as an Objective Function
Without loss of generality, consider a discrete BN. We formulate a backwards probability measure $P^{b}_{kl}$ for computing inference akin to the rearrangement of Equation (3), which forms the basis of inference techniques such as variable elimination [5].
@@
Define $\mathcal{F}()$ as a normalised count of the number of clock cycles associated with performing all the multiplications and additions, where each multiplication is assumed to have a normalised count of $1$ cycle, each addition $0.6$ and each division (i.e. ratio) $3$, based on Intel Skylake CPU latencies [6]. Hence,
@@ 5 Discussion
Clustering has been applied in a variety of approaches to improve the computational feasibility of BN learning, whether based on constraints, scoring (i.e. using a cost function) or both [10]. In this context, the vast majority of existing literature has focused on cost functions based on inclusion or exclusion of edges. Cost functions that vary with the mapping of nodes to different clusters are currently understudied. Such flexibility enables explicit optimisation of a given BN and its cluster mapping which could potentially lead to better BN learning and inference outcomes. However, finding globally optimal (i.e. least cost) cluster mappings in a BN incurs additional complexity in the form of cost dependency, where the local cost of a cluster is dependent on the mapping of connected nodes and the global cost is typically the sum of local costs.
Previous BN literature has used decomposability of scoring functions, such as through additive costs, to allow for local search procedures and/or optimisation of the score function; some even use dynamic programming and graph layers [7, 12, 18]. Our work builds upon the application of dynamic programming and layers in BNs to facilitate global optimisation with arbitrary non-negative cost functions and cost dependency. This latter property has not been explicitly studied in the existing BN literature.
The proposed algorithm, DCMAP, is proven to find all of the least cost cluster mapping solutions through conditional optimality of layers and branching (Section 3.2.3). In addition, it is theoretically shown that an equally informed algorithm cannot find these solutions in fewer iterations. This is important as optimal clustering is NP-hard in general [4] and cost dependency would exacerbate this complexity. For instance, a breadth first search on the same one time-slice case study is unable to find any optimal solution even with ten times more iterations than DCMAP. This arises even though the number of iterations needed is similar to DCMAP for the simple illustrative BN (Fig. 1).
@@ 5.1 Case Study Discussion
We developed a case study of BN clustering where clusters were used as a means to compute (i.e. infer) the marginal probabilities of a complex systems seagrass DBN. By explicitly defining the cost function as the computational cost of inference given a cluster mapping, a speed-up of 22.8 and 426 times was achieved for a one and two time-slice version of the DBN, respectively. In itself, a substantial reduction of computational complexity through clustering is useful as BN inference plays a key role in prediction and model learning, and methods for that such as Markov Chain Monte Carlo (MCMC) [8] can require thousands of iterations. MCMC can also help better quantify uncertainty for parameter and structure learning of BNs [5] and DBNs [24] through the Bayesian framework. Explicit optimisation of BN clusters using DCMAP could be a key enabler for future research in BN learning and updating, especially for complex systems DBNs like our seagrass case study.
@@
Although our case study was for a discrete BN with between two and fourteen states per node with corresponding CPTs, the proposed DCMAP algorithm equally applies to discrete or continuous or hybrid BNs, DBNs, Object Oriented Bayesian Networks (OOBNs) [13] and Markov models in general, all of which can be represented as a DAG; the key adaptation required is the development of an application-specific cost function. Note that dynamic and templated models alike can be unrolled [26] into a BN and cluster mapping performed on the result. In practice, for a $k^{th}$ order dynamic model, running cluster mapping on $k$ slices is sufficient to find optimal clusters that can then be replicated over all slices.
@@
We have theoretically shown that an equally informed algorithm cannot find all optimal solutions in fewer iterations, further study is needed to better understand the computational complexity of DCMAP and how it might vary with different BN graph structures and cost functions. Early results from the case study suggest a complexity of $O(n^{2})$ iterations to find the first least cost cluster mapping and $O(an^{2})$ iterations for all optimal solutions, where $n$ is the number of nodes and $a$ some positive integer. This aligns with previous findings of $O(n^{2})$ complexity for clustering in the absence of cost dependency [29]. However, the computational complexity could potentially vary significantly depending on the cost function employed and also the structure of the network (e.g. dense versus sparse). This is an avenue for future research.
Despite the theoretical optimality of DCMAP, scalability in practice is a challenge for large networks with thousands of nodes. Potentially, the template behind a DBN or OOBN like the two-slice DBN case study, and/or near-optimality could be investigated as approaches for large networks. Better understanding of the convergence characteristics and early termination of clustering, including in the impact of cost functions and network structures, is needed. Additionally, even though proposed cluster mappings are always unique (Lemma 3.3), label switching is an area for future study to better understand the extent of the issue and strategies for mitigation if needed.
@@ 5.2 Future Work: Entropy
In light of the generalisability of DCMAP given its ability to optimise for arbitrary, non-negative and dependent cost functions, one key area of future work is in the development of objective functions for specific usage scenarios. One such possibility is entropy, which for BNs is defined generally as the sum of the entropy associated with each scenario over all possible scenarios, where a scenario might be defined in terms of one or more nodes [23, 20]. Consider a discrete BN. The entropy associated with evidence for a given scenario indexed by $k$ is:
@@
Consider a special case of clustering where specific evidence scenarios are to be applied at one or more nodes and at one or more points in time in a DBN, with the goal of minimising global entropy. Thus, each node can be assigned to one of two clusters: (cluster 1) observed using virtual evidence $\delta(X^{\prime})$ [22] where each observation is independent, or (cluster 2) not observed. Note that virtual evidence provides greater flexibility in modelling uncertainty compared to hard evidence where a state is observed or not. Dependent cluster mapping is necessary to optimise over all possible combinations of observed nodes and their interdependencies, which arise from propagation of evidence for inferring marginal probabilities for each scenario $p_{ijk}$ (20). DCMAP can approximate this marginal probability $p_{ijk}$ for every node $X_{i}$ via backwards propagation with evidence to layer $l$ and no evidence from forward propagation:
@@
Clusters have been used to improve the computational efficiency of inference, and learn conditional probabilities and BN structures [5, 18]; potentially, optimal clusters can further enhance these tasks. In addition, using cost functions such as entropy, clusters could also have a practical interpretation. In non-homogeneous Hidden Markov Models or DBNs for instance, clusters could be interpreted as homogeneous portions of a non-homogeneous system, akin to changepoint detection or model switching [24, Boys2000]. Potentially, optimisation of clusters with respect to entropy could contribute to new methods in BN learning and prediction, and is a promising area for future research that is enabled by DCMAP.
[…]
Supplementary Information
6 Supplementary Information: Super-additive Objective Functions